# The Role of Electric Grid Research in Addressing Climate Change

### v1

@@ Abstract
The electricity sector plays two pivotal roles in tackling climate change. First, according to the recent IPCC report [1], as of 2019, the energy sector contributes to approximately 34% of global greenhouse gas emissions. Therefore, cleaning up the electricity sector itself is a major step towards the goal of reducing gigaton-level carbon emissions for the entire planet by 2050. Second, for many carbon management or reduction technologies, the best route to achieve a speedy and scalable impact is through large-scale integration into the electric grid.
@@
Achieving gigaton-level carbon emission reduction through the power grid necessitates the expansion of the electrical systems, which is now more intertwined with weather and climate than ever before (Figure 1). For example, due to their variable nature of renewable resources, integrating them into the power grid significantly impacts planning processes, which used to operate under the generation-following-demand paradigm [10]. The emergence of long-term energy storage [11] and demand response technologies [12, 13, 14] has led to the concept of demand-following-generation as an added mechanism to achieve overall power balancing during operation. However, the necessary long-term storage for the power grid, e.g., for ‘dunkelflaute’ events [15], can only be accurately planned through climate model simulations [16]. In this article, we ask two key questions. First, what power grid researchers should prioritize as we integrate more and more carbon-neutral technologies? Second, how do we effectively integrate climate research with power grid research to address the emerging complexities? The following two sections provide an answer to these two questions.
Figure 1: The crucial role of the electric grid research in tackling climate change. This schematic illustrates how electric grid research interacts with climate change research. ‘Key Role 1’ represents the decarbonization of the power generation sector, while ‘Key Role 2’ denotes the integration of other sectors—such as transportation, industrial heating/cooling, computing loads, and household appliances—into the electric grid, facilitating their decarbonization. We show that climate and weather models must be utilized for power grid operations and planning, with the power grid providing essential feedback to enhance the climate models. Finally, a climate-aware market redesign and policy support are essential to leveraging the power grid in combating climate change.
@@
Power grids around the world evolved to cater to regional energy demand, supported by local policy mandates and unique generation characteristics (e.g., the Pacific Northwest in the United States has plenty of hydro-electric potential [43], western Texas has an abundance of wind resources [44]). Global renewable energy potential is not uniform, and the distribution can be impacted by climate change [21]. While climate change is a planetary-scale problem, each region faces its unique climate-related challenges (e.g., the wildfires in California [45] and hurricanes on the Gulf and east coast of the United States [46]). Therefore, the interdependence between energy and climate systems can result in regional energy security concerns with the growing prominence of renewable energy sources, and power system planners need to be cognizant of this relationship while planning.
@@
Figure 2: Power grid resource adequacy assessment utilizing climate data. The climate data, obtained from both high-resolution (0.25 degrees) and low-resolution (1 degree) global climate simulations, include atmospheric temperature, relative humidity, dew point, wind speed, and solar radiation flux. These datasets enable more accurate predictions of renewable generation and load demands for the power grid, e.g., from 2033 to 2043. They enhance traditional resource adequacy frameworks that already include load growth models, generator capacity expansion models, and generator planned outage models, which are essential for determining effective system-wide loads and generations. By integrating these comprehensive climate datasets with traditional models, we can better assess resource adequacy and identify potential reliability issues under various climate scenarios.
@@
Enabled by improved climate models, we highlight three innovative market designs to improve resource adequacy issues in deregulated environments. Firstly, the capacity market is one way of ensuring that all generators will be present to participate in the spot market. Regarding renewable generation, Joskow [36] discussed a hybrid framework for capacity and spot markets, where long-term power purchase agreements with wind, solar, and storage developers are competitively procured, which the author called the “competition for the market,” and incorporating it with “competition in the market” through short-term energy markets. Access to a good climate model would raise confidence among renewable energy developers to develop more renewable energy farms and participate in the market without needing additional support. Secondly, enabled by technological innovations on the consumer end and improved accuracy of consumption patterns, we frequently see electricity demand side participation in the real world, e.g., load resources participation in Texas electricity market [12], energy coupon [13], and direct load control [14]. There has been an increasing demand for an edge-based market that allows consumers to trade their excess energy and integrate it with the wholesale energy market. Thirdly, as we integrate more and more renewable power plants and continue to electrify other sectors, we may run out of transmission line capacity. Innovative solutions with storage (e.g., deploying storage devices at both ends of the transmission lines) enabled by data analytics and climate models would facilitate optimal utilization of transmission resources. Alternatively, energy efficiency [63] is another solution to reduce system-wide demand itself.
@@ Conclusion
Acknowledgments The work of L.X., S.M., Q.Z. and P.C. is supported in part by Texas A&M Energy Institute, College of Arts and Sciences at Texas A&M University, and Texas A&M Engineering Experiment Station.
@@
Author contributions L.X., S.M. T.H. and Q.Z. conceived and designed the paper. P.C. and D.J.H. contributed material and analysis tools. L.X., S.M. T.H., and Q.Z. drafted the paper with input from all co-authors. All authors read and approved the final version of the paper.
@@
[…]
Peer review information The authors sincerely thank all the anonymous reviewers for significantly improving the quality of the paper.

### v2

@@ Abstract
The electricity sector plays two pivotal roles in tackling climate change. First, according to the recent IPCC report [1], as of 2019, the electricity and heating sector contributes to approximately 23% of global greenhouse gas emissions. Therefore, cleaning up the electricity sector itself is a major step towards the goal of reducing gigaton-level carbon emissions for the entire planet by 2050. Second, for many carbon management or reduction technologies, the best route to achieve a speedy and scalable impact is through large-scale integration into the electric grid.
@@
Therefore, achieving gigaton-level carbon emission reduction through the power grid necessitates the expansion of the electrical systems, which is now more intertwined with weather and climate than ever before (Figure 1). For example, due to their variable nature of renewable resources, integrating them into the power grid significantly impacts planning processes, which used to operate under the generation-following-demand paradigm [10]. The emergence of long-term energy storage [11] and demand response technologies [12, 13, 14] has led to the concept of demand-following-generation as an added mechanism to achieve overall power balancing during operation. However, the necessary long-term storage for the power grid, e.g., for ‘dunkelflaute’ events [15], can only be accurately planned through climate model simulations [16]. In this article, we ask two key questions. First, what power grid researchers should prioritize as we integrate more and more carbon-neutral technologies? Second, how do we effectively integrate climate research with power grid research to address the emerging complexities? The following two sections provide an answer to these two questions.
Figure 1: The crucial role of the electric grid research in tackling climate change. This schematic illustrates how electric grid research interacts with climate change research. ‘Key Role 1’ represents the decarbonization of the power generation sector, while ‘Key Role 2’ denotes the integration of other sectors—such as transportation, industrial heating/cooling, computing loads, and household appliances—into the electric grid, facilitating their decarbonization. We show that climate and weather models must be utilized for power grid operations and planning, with the power grid providing essential feedback to enhance the climate models. Finally, a climate-aware market redesign and policy support are essential to leveraging the power grid in combating climate change. The icons used in this figure are sourced from flaticon.com.
@@
First, power grids around the world evolved to cater to regional energy demand, supported by local policy mandates and unique generation characteristics (e.g., the Pacific Northwest in the United States has plenty of hydro-electric potential [43], western Texas has an abundance of wind resources [44]). Second, global renewable energy potential is not uniform, and the distribution can be impacted by climate change [21]. Third, while climate change is a planetary-scale problem, each region faces its unique climate-related challenges (e.g., the wildfires in California [45] and hurricanes on the Gulf and east coast of the United States [46]). Therefore, the interdependence between energy and climate systems can result in regional energy security concerns with the growing prominence of renewable energy sources, and power system planners need to be cognizant of this relationship while planning.
@@
Figure 2: Power grid resource adequacy assessment utilizing climate data. The climate data, obtained from both high-resolution (0.25 degrees) and low-resolution (1 degree) global climate simulations, include atmospheric temperature, relative humidity, dew point, wind speed, and solar radiation flux. These datasets enable more accurate predictions of renewable generation and load demands for the power grid, e.g., from 2033 to 2043. They enhance traditional resource adequacy frameworks that already include load growth models, generator capacity expansion models, and generator planned outage models, which are essential for determining effective system-wide loads and generations. By integrating these comprehensive climate datasets with traditional models, we can better assess resource adequacy and identify potential reliability issues under various climate scenarios. Figure adapted from ref. [47]. The icons used in this figure are sourced from flaticon.com.
@@
We highlight three innovative market designs to improve resource adequacy issues in deregulated environments, which could be enabled by improved climate models. Firstly, the capacity market is one way of ensuring that all generators will be present to participate in the spot market. Regarding renewable generation participation in capacity markets, Joskow [36] discussed a hybrid framework for capacity and spot markets, where long-term power purchase agreements with wind, solar, and storage developers are competitively procured, which the author called the “competition for the market,” and incorporating it with “competition in the market” through short-term energy markets. Access to a good climate model would raise confidence among renewable energy developers to develop more renewable energy farms and participate in the market without needing additional support. Secondly, enabled by technological innovations on the consumer end and improved accuracy of consumption patterns, we could frequently see electricity demand side participation in the real world, e.g., load resources participation in Texas electricity market [12], energy coupon [13], and direct load control [14]. There has been an increasing demand for an edge-based market that allows consumers to trade their excess energy and integrate it with the wholesale energy market. Thirdly, as we integrate more and more renewable power plants and continue to electrify other sectors, we may run out of transmission line capacity. Innovative solutions with storage (e.g., deploying storage devices at both ends of the transmission lines) enabled by data analytics and climate models would facilitate optimal utilization of transmission resources. Alternatively, energy efficiency [63] is another solution to reduce system-wide demand itself.
@@ Conclusion
Acknowledgments The work of L.X., S.M., Q.Z. and P.C. is supported in part by Texas A&M Energy Institute, College of Arts and Sciences at Texas A&M University, and Texas A&M Engineering Experiment Station. The work of T.H. is supported by US National Science Foundation Grant 2328205. The work of L.X. is also supported in part by US National Science Foundation Grant ECCS-2038963.
@@
Author contributions L.X., S.M. T.H. and Q.Z. conceived and designed the paper. P.C. and D.J.H. contributed material and analysis tools. L.X., S.M. T.H., and Q.Z. drafted the paper with input from all co-authors. L.X., S.M. T.H., Q.Z., P.C., D.J.H., and M.S. read and approved the final version of the paper.